OpenStudy (anonymous):

Verify the identity. cos 4x + cos 2x = 2 - 2 sin^2 2x - 2 sin^2 x

OpenStudy (anonymous):

@jim_thompson5910 @seiga

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

...This is relatively straightforward Let's work on the RHS . We will use the fact that cos(A+B) = cosAcosB - sinAsinB Now cos(4x) = cos(2x + 2x) = cos(2x)cos(2x) - sin(2x)sin(2x) = cos^2(2x) - sin^2(2x) = LHS, as required